FIND THE GENERAL SOLUTION

Need ALL 4. Differential Equations class

9. (t) = ( ) (1) 10. 7(t) = 2 11. x(t) : | ar(t) 12. () =( ) (t
Answer #1

9.

Given x = AX, 3 -27 |X, X= 4 -1 x = 3x – 2y → (1) v = 4x- y → (2) (1) and (2) in operator form (D-3)x+2y = 0 → (3) where D